OpenStudy (1dfanatic):

You have a bag containing three red, seven green, and six blue pens. You choose tow pens at random. Find P(blue and blue) with replacing. Make sure to show/explain your work.

OpenStudy (anonymous):

so let's see there are 16 pens.

OpenStudy (anonymous):

Choosing blue is 3/8 chance

OpenStudy (anonymous):

second time is 1/3

OpenStudy (anonymous):

multiply to find chance

OpenStudy (welshfella):

first find probability of drawing a blue on a single draw. since the 2 draws are independent the above probability is multiplied by itself to give the answer you want.

OpenStudy (anonymous):

The probability of picking two blue pens is approx 12.5%

OpenStudy (1dfanatic):

i got 1/3

OpenStudy (welshfella):

No the probabiluty of a second blue is also 3/8 because the ball is replaced.

OpenStudy (welshfella):

so P(blue and blue) = 3/8 * 3/8

OpenStudy (1dfanatic):

then it'd be 9/64

OpenStudy (welshfella):

yes

OpenStudy (1dfanatic):

thank you

OpenStudy (welshfella):

you understand why 1 draw = 3/8? 6 out of 16 = 3/8

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

dahmanman would have been right if the ball was not replaced

OpenStudy (welshfella):

- because there would have been 5 blues out of a total of 15 balls, which = 1/3